OpenStudy (anonymous):

find the slope of the line that passes through each pair of points. (7/3,4/3),(-1/3,2/3)

OpenStudy (lgbasallote):

do you have an initial answer?

OpenStudy (lgbasallote):

i mean if you have tried solving it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

before you helped me I was trying to solve by graphing each line & it was taking forever. I turned the fractions into decimals and got m=1/2 at first but that sounded incorrect. Now I dont know how to solve

OpenStudy (lgbasallote):

yup it's not 1/2

OpenStudy (lgbasallote):

first step is 2/3 - 4/3 remember in subtracting fractions you just subtract the numeratirs and copy the denominator so it becomes \[\frac{2-4}[3}\] got it?

OpenStudy (lgbasallote):

i mean \[\frac{2-4}{3}\]

OpenStudy (anonymous):

oh.... I did not know that. I'm going back to high school after 10 yrs. So I do not remember that. ..... Ok so if it's the same denominator we use for both?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

so \[\LARGE slope = \frac{\frac{-1 - 7}{3}}{\frac{2-4}{3}}\]

OpenStudy (anonymous):

so then we have \[\left(\begin{matrix}-2/3 \\ -8/3\end{matrix}\right)\]

OpenStudy (lgbasallote):

oops yeah

OpenStudy (lgbasallote):

sorry i got mine topsy turvy

OpenStudy (lgbasallote):

\[\LARGE slope = \frac{\frac{-8}{3}}{\frac{-2}{3}}\] now...the 3 cancels out

OpenStudy (lgbasallote):

so you're left with \[\frac{-8}{-2}\]

OpenStudy (lgbasallote):

because \[\LARGE \frac{\frac{a}{c}}{\frac{b}{c}} = \frac{a}{b}\]

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

yes I got it. Does it turn into a positive since both numerators are negative?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

ok. Great thank you. Now that I have that part and gave u a medal. I'm going to post another question since I do not get the next section.